[(x-1)3]8 the 3 and the 8 are exponents
Answer by checkley71(8403)

((X-1)^3)^8=(X-1)^24
THE FORMULA FOR EXPANDING THIS TERM IS:
X^24-(24/1)X^23+[(24*23)/(1*2)]X^22-[(24*23*22)/(1*2*3)]X^21+[(24*23*22*21)/(1*2*3*4)]X^20-[(24*23*22*21*20)/(1*2*3*4*5)]X^19+[(24*23*22*21*20*19)/(1*2*3*4*5*6)]X^18 etc.
EXPANDED WE HAVE:
X^24-24X^23+276X^22-2,024X^21+10,626X^20-42,504X^19+134,596X^18 etc.
---------------------------------------------------
NOTE THE CHANGE OF SIGNS + - + - + - + - + - ETC.
THERE WILL BE A TOTAL OF 25 TERMS WITH THE LAST TERM=1^24=1.